# TOGLL: Correct and Strong Test Oracle Generation with LLMs

### v1

@@ Abstract
Test oracles play a crucial role in software testing, enabling effective bug detection. Despite initial promise, neural-based methods for automated test oracle generation often result in a large number of false positives and weaker test oracles. While LLMs have demonstrated impressive effectiveness in various software engineering tasks, including code generation, test case creation, and bug fixing, there remains a notable absence of large-scale studies exploring their effectiveness in test oracle generation. The question of whether LLMs can address the challenges in effective oracle generation is both compelling and requires thorough investigation.
In this research, we present the first comprehensive study to investigate the capabilities of LLMs in generating correct, diverse, and strong test oracles capable of effectively identifying a large number of unique bugs. To this end, we fine-tuned seven code LLMs using six distinct prompts on the SF110 dataset. Utilizing the most effective fine-tuned LLM and prompt pair, we introduce TOGLL, a novel LLM-based method for test oracle generation. To investigate the generalizability of TOGLL, we conduct studies on 25 large-scale Java projects. Besides assessing the correctness, we also assess the diversity and strength of the generated oracles. We compare the results against EvoSuite and the state-of-the-art neural method, TOGA. Our findings reveal that TOGLL can produce 3.8 times more correct assertion oracles and 4.9 times more exception oracles. Moreover, our findings demonstrate that TOGLL is capable of generating significantly diverse test oracles. It can detect 1,023 unique bugs that EvoSuite cannot, which is ten times more than what the previous SOTA neural-based method, TOGA, can detect.
@@ I Introduction
The efficacy of test oracles is determined by their ability to detect bugs. For effective bug detection, test oracles should be correct and strong. A test oracle is considered correct if it aligns with the expected program behavior, thereby avoiding false alarms. Additionally, a test oracle is deemed strong if it can detect deviations from intended program behavior. Just because a test oracle is correct does not mean it is strong, as demonstrated in  [16]. Even though developer-written test oracles are proven more stronger than automated test oracles [51], manual generation of test oracles is both costly and time-consuming [18, 3]. To address this, researchers have explored methods to automate oracle generation.
Figure 1: Overview of our approach to explore LLM-based oracle generation and to evaluate TOGLL.
@@
Large language models (LLMs) have shown promise in software engineering tasks such as code completion [25], program comprehension [29], automated repair [19], and test generation [42, 24, 43]. Most prior work on ML-based test generation has focused on the generation of test prefixes – inputs to the system under test — with the goal of yielding substantial test coverage. CodaMosa [24] uses LLMs to escape coverage plateaus in search based test input generation and find it can improve coverage on small to moderate Python programs. A recent attempt to use LLMs to generate test prefixes for large-scale Java programs was not very successful [43]; it achieved only 2% statement coverage on large-scale Java projects compared to over 90% for EvoSuite’s [12] search-based approach. The LLM-based TESTPILOT [42] approach generates both test prefixes and test oracles for JavaScript. It produces better coverage than state-of-the-art search-based JavaScript test generation techniques on a variety of small programs. With regard to test oracles, TESTPILOT was shown to generate significant numbers of non-trivial test assertions that are distinct from those generated by other methods, but this work did not evaluate the strength of the generated test oracles for fault detection.
@@
In this paper, rather than task the LLM with generating both test prefix and test oracle, as TESTPILOT did, we decouple these problems. We utilize state-of-the-art test generation methods, such as EvoSuite, to create test prefixes and concentrate on leveraging LLMs for test oracle generation. Initially, we fine-tune seven code LLMs, with sizes ranging from 110M to 2.7B parameters, employing six prompt formats that vary in the information they incorporate about the method under test (MUT) and the test prefix. Employing the most effective model and prompt pair, we introduce our LLM-based method, TOGLL. To evaluate TOGLL’s generalizability on unseen projects, we conduct a comprehensive study on 25 real-world Java projects and compare the correctness of the generated oracles with those produced by the state-of-the-art neural method TOGA [10] and EvoSuite. Subsequently, we assess the strength of the generated oracles through a large-scale bug detection study, highlighting the unique advantages of LLM-generated oracles over alternative methods. To the best of our knowledge, this represents the first extensive study to explore the application of LLMs in test oracle generation, examining various code LLMs, prompt formats, and two large datasets, while evaluating multiple aspects of the generated oracles, including correctness, diversity, and effectiveness in bug detection.
@@
  • Evaluating the generalizability of fine-tuned LLMs in generating effective test oracles for unseen projects and compare them with SOTA methods.
  • Evaluating the impact of different contextual information on oracle generation accuracy through fix specially designed prompts.
@@ II Approach
Figure 1 presents an overview of our research approach. Step 1 of Figure 1 depicts fine-tuning a collection of LLMs with different prompting approaches for test oracle generation task; we discuss this in Section II-A. Step 2 of Figure 1 selects the best performing fine-tuned model and prompt from step 1 to define TOGLL and to evaluate its performance on unseen data. Step 3 assesses the correctness of TOGLL generated oracles through test execution; we discuss this in Section II-B1. Finally, step 4 assesses the strength of TOGLL generated oracles using mutation testing; we discuss this in Section II-C.
@@ II-A Supervised Fine-tuning
For supervised fine-tuning, we use the SF110 dataset, construct six different prompts, and fine-tune seven different code LLMs.
II-A1 SF110 Training Data
We construct the training dataset using the SF110 EvoSuite benchmark [13], which comprises 110 open-source Java projects with over 23,000 Java classes sourced from the SourceForge repository. This benchmark is widely recognized for unit test generation. To construct our dataset, we utilize the test cases generated by EvoSuite for each project. Given that a single test case can contain multiple assertion oracles or even a combination of assertion and exception oracles, we processed each test case. This involved decomposing them into individual test cases, ensuring that each contained only a single assertion or exception oracle. For each decomposed test case, we extracted the method under test along with its associated documentation to prepare the dataset. The dataset is comprised of tuples $((p_{i},m_{i},d_{i}),o_{i})$, where, $p_{i}$ is the test prefix, $m_{i}$ is the method under test, and $d_{i}$ is the docstring (if available) and $o_{i}$ is the ground truth test oracle, either assertion or exception. The dataset has been partitioned into three distinct subsets: 80% for training, 10% for testing, and another 10% for validation.
@@ II-A2 Prompt Designing
Even though LLMs are powerful, they are not specifically designed and pre-trained for test oracle generation. Thus, fine-tuning with effective prompts is necessary for optimal performance. We have designed six different prompts that incorporate increasing amounts of context to investigate their effectiveness in oracle generation.
  • Prompt 1 (P1) consists of only the test prefix (‘prefix’). This prompt is suitable when addition details on the MUT or documents are not available.
  • Prompt 2 (P2) additionally includes the method documentation strings (‘prefix + [sep] + doc’) . The ‘[sep]’ separator token is different based on different LLM and their respective tokenizer.
  • Prompt 3 (P3) instead of docstrings, includes the method signature (‘prefix + [sep] + mutsig’). This prompt is suitable when neither the entire method implementation nor docstring available. Our experimental study in RQ1 shows that simply adding method signature can significantly improve accuracy.
  • Prompt 4 (P4) includes both method signature and docstrings in addition to prefix (‘prefix + [sep] + doc +[sep] + mutsig’).
  • Prompt 5 (P5) includes the code for the entire method under test (‘prefix + [sep] + mut‘).
  • Prompt 6 (P6) adds the docstring to P5 (‘prefix + [sep] + doc + [sep] + mut‘).
@@ II-A3 Code LLMs
In our study, we use pre-trained auto-regressive decoder models that are widely used, available on Hugging Face, and compatible with our GPU resources.
@@ II-B Unseen Inference
II-B1 Unseen Inference Data
To investigate the generalizability of the fine-tuned model, we utilized the dataset described in  [16], comprising 25 large-scale industrial standard Java projects. Of these, 17 originate from the Apache Commons Proper [38], while the remainder are sourced from GitHub. These latter projects are characterized by large codebases with multiple modules, with up to 10k active users per project, and are frequently employed in software engineering research for the evaluation of test cases and test oracles  [20, 41, 51]. Given their broad coverage of 21 domains and significant variation in developer metrics, the 25 artifacts selected for this study provide a robust framework for assessing fine-tuned model’s ability to generalize across a spectrum of real-world projects. Overall, the dataset consists of 271K source lines of code (SLOC), 331K lines of JavaDoc, 1214K test suite SLOC and 223.5K test cases. Similar to the SF110 dataset, each input sample of this dataset contains the test prefix, MUT and docstring.
@@ II-B2 TOGLL
Based on the fine-tuning results, we selected the top two performing models (CodeGen-350M, CodeParrot-110M) based on their accuracy and also chose the three best-performing prompts (numbers P4, P5, and P6). We then conducted a small-scale study, performing unseen inference with these two models and three prompts on a few projects from the unseen dataset discussed in Section II-B1. Our results suggest that even though CodeGen-350M is three times larger than CodeParrot-110M, CodeParrot-110M fine-tuned with P6 performs better on the unseen project than CodeGen on several metrics: total input processed, compilation error, and false positive count. Furthermore, the inference time with CodeGen-350M is longer than CodeParrot. Consequently, we define TOGLL as a fine-tuned CodeParrot-110M operating on P6. We perform rest of the study with TOGLL on the unseen dataset II-B1.
II-B3 Accuracy Computation Through Test Validation
With TOGLL, we generate test oracles for all 25 projects from the unseen dataset. This dataset framework includes all necessary artifacts for executing the generated test oracles when integrated with the test prefixes. Consequently, we integrate the generated oracles into the test cases and execute the complete test suites. We calculate accuracy as the success rate, which denotes the percentage of non-empty test oracles that were successfully passed during test execution, indicating their alignment with expected program behavior. We compute the success rate using Equation 1
$\textrm{SuccessRate}(A)=\frac{T-(T_{ce}+T_{fp}+T_{em})}{T}$ (1)
where $T$ is the number of total test prefixes in the artifact $A$, $T_{ce}$ is the number of compilation errors, $T_{fp}$ is the number false positive tests, and $T_{em}$ the number of empty oracles. We assess the success rates for both assertion and exception oracle predicted by TOGLL and compare these rates to those of our baseline, TOGA, as discussed in Section II-B4.
II-B4 Baseline (TOGA)
@@
TOGA is a neural method for both assertion and exception oracle generation [10]. TOGA utilizes CodeBERT [11] model as backbone and fine-tuned for the oracle generation task. TOGA takes the test prefix generated by EvoSuite, its associated MUT and docstring (if available) and generates either an assertion oracle or predicts an exception oracle. An example of EvoSuite generated test cases with assertion and exception oracle is shown in Figure 2. The prefix part is marked with yellow color. The first test prefix involves pushing the number 10 onto the stack, followed by popping this value from the stack. For this scenario, an assertion oracle is anticipated to validate that the value retrieved is indeed 10. The second test attempts to pop from an empty stack, expecting the MUT to throw an exception. If no exception is thrown, the test should fail.
@@
We selected TOGA as our baseline method for several reasons: 1) TOGA represents the SOTA neural method, having surpassed specification, search, and neural-based techniques by detecting 57 bugs in Defects4J [20]; 2) TOGA utilizes the same set of information as P6, which is our most comprehensive prompt, thereby providing a fair basis for comparison; 3) TOGA has been extensively evaluated on the same unseen dataset that we also employ in our study, ensuring consistency in our comparative analysis [16].
@@
Figure 2: EvoSuite-Generated Test Cases with Assertion and Exception Oracles. The prefix part is marked with yellow color.
II-C Mutation Testing
@@
To assess the strength of the generated oracles, we perform mutation testing that introduces small changes, or mutations, to the source code to create numerous slightly altered versions of the software, called mutants[8, 35]. Studies showed that there is a statistically significant correlation between mutant detection and real fault detection [21, 1, 34]. A test suite’s ability to detect these mutants can be a direct measure of its bug detection effectiveness [51, 41]. We conduct mutation testing on the cases where the assertion generated by TOGLL could be executed, meaning the assertions were neither syntactically incorrect nor false positives. From the unseen dataset, we consider the same set of test cases with their test oracles to enable comparison. For each of the 25 unseen artifacts, we perform mutation testing and record the total number of mutants detected by each approach. Additionally, we record the total number of unique mutants which indicates the unique capability of each oracle generation method. Besides comparing with EvoSuite, we also compare the results with TOGA.
III Experimental Study
To assess the efficacy of LLM-generated test oracles, this study investigates several key areas: effectiveness of different code LLMs in generating test oracles, the impact of different prompts on the oracle generation accuracy, the ability of fine-tuned models to generalize to unseen data, comparative accuracy among different oracle types, diversity of the generated oracle and the overall efficacy of these oracles in detecting software bugs. To this end, we answer the following research questions.
RQ4: How strong are LLM-generated assertions at identifying unique bugs? TOGLL generated oracles kill nearly 10 times more unique mutants than prior SOTA method and fall only 17% short of EvoSuite’s bug detection effectiveness.
@@ III Experimental Study
TABLE I: Test Oracle Generation Performance of Code LLMs for Different Prompts. Top two values for each prompt are shown in bold font. The lower row shows prompt-wise average.

### v2

@@ Abstract
Test oracles play a crucial role in software testing, enabling effective bug detection. Despite initial promise, neural methods for automated test oracle generation often result in a large number of false positives and weaker test oracles. While LLMs have shown impressive effectiveness in various software engineering tasks, including code generation, test case creation, and bug fixing, there remains a notable absence of large-scale studies exploring their effectiveness in test oracle generation. The question of whether LLMs can address the challenges in effective oracle generation is both compelling and requires thorough investigation.
In this research, we present the first comprehensive study to investigate the capabilities of LLMs in generating correct, diverse, and strong test oracles capable of effectively identifying a large number of unique bugs. To this end, we fine-tuned seven code LLMs using six distinct prompts on a large dataset consisting of 110 Java projects. Utilizing the most effective fine-tuned LLM and prompt pair, we introduce TOGLL, a novel LLM-based method for test oracle generation. To investigate the generalizability of TOGLL, we conduct studies on 25 unseen large-scale Java projects. Besides assessing the correctness, we also assess the diversity and strength of the generated oracles. We compare the results against EvoSuite and the state-of-the-art neural method, TOGA. Our findings reveal that TOGLL can produce 3.8 times more correct assertion oracles and 4.9 times more exception oracles than TOGA. Regarding bug detection effectiveness, TOGLL can detect 1,023 unique mutants that EvoSuite cannot, which is ten times more than what TOGA can detect. Additionally, TOGLL significantly outperforms TOGA in detecting real bugs from the Defects4J dataset.
@@ I Introduction
The efficacy of test oracles is determined by their ability to detect bugs. For effective bug detection, test oracles should be correct and strong. A test oracle is considered correct if it aligns with the expected program behavior, thereby avoiding false alarms. Additionally, a test oracle is deemed strong if it can detect deviations from intended program behavior. Just because a test oracle is correct does not mean it is strong, as demonstrated in  [5].
Figure 1: A general overview of our research approach to investigating LLM-based automated oracle generation.
@@
Large language models (LLMs) have shown promise in automating diverse software engineering tasks such as code completion [18], program comprehension [19], automated repair [20], and test generation [21, 22, 23]. Most prior work on ML-based test generation has focused on the generation of test prefixes – inputs to the system under test — with the goal of yielding substantial test coverage. CodaMosa [22] uses LLMs to escape coverage plateaus in search based test input generation and find it can improve coverage on small to moderate Python programs. A recent attempt to use LLMs to generate test prefixes for large-scale Java programs was not very successful [23]; it achieved only 2% statement coverage on large-scale Java projects compared to over 90% for EvoSuite’s [24] search-based approach. The LLM-based TESTPILOT [21] approach generates both test prefixes and test oracles for JavaScript. It produces better coverage than state-of-the-art search-based JavaScript test generation technique on a variety of small programs. With regard to test oracles, TESTPILOT was shown to generate significant numbers of non-trivial assertions that are distinct from those generated by other methods, but this work is limited to JavaScript and did not evaluate oracles correctness, diversity and strength aspects thoroughly.
@@
In this paper, rather than task the LLM with generating both test prefix and test oracle, as TESTPILOT did, we decouple these problems. We utilize state-of-the-art test generation methods, such as EvoSuite, to create test prefixes and concentrate on leveraging LLMs for test oracle generation. Initially, we fine-tune seven code LLMs, with sizes ranging from 110M to 2.7B parameters, employing six prompt formats that vary in the information they incorporate about the method under test (MUT), its documentation, and the test prefix. Employing the most effective model and prompt pair, we introduce our LLM-based method, TOGLL.
It is important to note that TOGLL generated oracles are not regression oracles. Regression oracles are generated based on the executed program behavior. For example, EvoSuite executes the code, collects all actual results values, knows the types, and then generates oracles based on those. This is why regression oracles are not capable of detecting bugs in current program versions. TOGLL, in contrast, does not execute the code and is not aware of the test execution results. It uses Javadoc documentation, MUT/MUT signature and test prefix to generate oracles. Some of the prompts (see Table I) do not even use MUT (P1, P2) at all and some (P3, P4) only use MUT signature, not the entire implementation.
To evaluate TOGLL’s generalizability on unseen data, we conduct a comprehensive study on 25 real-world large-scale Java projects and compare the correctness of the generated oracles with those produced by the SOTA neural method, TOGA [17] and EvoSuite. Subsequently, we assess the strength of the generated oracles through a large-scale mutant detection study and a real bug detection study, highlighting the unique advantages of TOGLL-generated oracles over alternative methods. To the best of our knowledge, this represents the first extensive study to explore the application of LLMs in test oracle generation, examining seven code LLMs, six prompts, and three large datasets, while evaluating multiple aspects of the generated oracles, including correctness, diversity, and bug detection effectiveness.
@@
  • Evaluating the generalizability of fine-tuned LLMs in generating effective test oracles on 25 unseen projects and compare them with SOTA methods.
  • Evaluating the impact of different contextual information on oracle generation accuracy through six specially designed prompts.
@@ II Approach
Figure 1 presents an overview of our research approach. Step 1 of Figure 1 depicts fine-tuning a collection of LLMs with different prompts for test oracle generation task; we discuss this in Section II-A. Step 2 of Figure 1selects the best performing fine-tuned model and prompt from Step 1 to define TOGLL and perform oracle inference on the unseen data. Step 3 assesses the correctness of TOGLL generated oracles through test execution; we discuss this in Section II-B1. Finally, step 4 assesses the strength of TOGLL-generated oracles using mutation testing and real bug detection from Defects4J; we discuss this in Section III-D and III-E.
@@ II-A Supervised Fine-tuning
This section covers the datasets, prompts, and LLMs that are fine-tuned for the test oracle generation task.
II-A1 SF110 Dataset
We construct this dataset using the SF110 EvoSuite benchmark [25], which comprises 110 open-source Java projects with over 23,000 Java classes sourced from the SourceForge repository. This benchmark is widely recognized for unit test generation. To construct our dataset, we utilize the test cases generated by EvoSuite for each project. Given that a single test case can contain multiple assertion oracles or even a combination of assertion and exception oracles, we process each test case. This involve decomposing them into individual test cases, ensuring that each contains only a single assertion or exception oracle. For each decomposed test case, we extracted the method under test (MUT) along with its associated documentation to prepare the dataset. The dataset is comprised of tuples $((p_{i},m_{i},d_{i}),o_{i})$, where, $p_{i}$ is the test prefix, $m_{i}$ is the MUT, and $d_{i}$ is the docstring (if available) and $o_{i}$ is the ground truth test oracle, either assertion or exception. This dataset is used to fine-tuning the LLMs and has been partitioned into three distinct subsets: 90% for training, 5% for validation, and another 5% for testing.
@@ II-A2 Prompt Designing
Even though LLMs are powerful, they are not specifically designed and pre-trained for test oracle generation. Thus, fine-tuning with effective prompts is necessary for optimal performance. We have designed six different prompts, each tailored to a specific use case scenario.
  • Prompt 1 (P1) consists of only the test prefix (‘prefix’).
  • Prompt 2 (P2) additionally includes the MUT documentation strings (‘prefix + [sep] + doc’) . The ‘[sep]’ separator token is different based on different LLMs and their respective tokenizer.
  • Prompt 3 (P3) instead of docstrings, includes the MUT signature (‘prefix + [sep] + mutsig’).
  • Prompt 4 (P4) includes both MUT signature and docstrings in addition to prefix (‘prefix + [sep] + doc +[sep] + mutsig’).
  • Prompt 5 (P5) includes the code for the entire MUT (‘prefix + [sep] + mut’).
  • Prompt 6 (P6) adds the docstring to P5 (‘prefix + [sep] + doc + [sep] + mut’).
These prompts are designed to cover diverse real use case scenarios. For example, when documents are not available, P1, P3, and P5 can be used. When the MUT implementation is not available, P1 - P4 can be used, and when it is available, P5 or P6 can be used. Designing these prompts allows us to assess their relative accuracies and the impact of different information on the oracle generation performance.
@@ II-A3 Code LLMs
Decoder-only LLMs are most suitable for code generation tasks [26]. Smaller models are cost-effective, produce a lower carbon footprint [27], and can be more easily fine-tuned with limited GPU resources. Therefore, in our study, we prefer decoder-only LLMs that are pre-trained on various code generation tasks within the software development life cycle (SDLC)[28, 26], smaller in size, publicly available on Hugging Face[29], and well-documented for fine-tuning. We chose seven LLMs because they are pre-trained on multiple programming languages while meeting all the above criteria.
@@ II-B Unseen Inference
II-B1 OracleEval25 Dataset
To investigate the generalizability of the fine-tuned models, we utilize the dataset described in  [5], comprising 25 large-scale industrial standard Java projects. Of these, 17 originate from the Apache Commons Proper [35], while the remainder are sourced from GitHub. These latter projects are characterized by large codebases with multiple modules, with up to 10k active users per project, and are frequently employed in software engineering research for the evaluation of test cases and test oracles  [36, 6, 7]. Given their broad coverage of 21 domains and significant variation in developer metrics, the 25 artifacts selected for this study provide a robust framework for assessing fine-tuned model’s ability to generalize across a spectrum of real-world projects. Overall, the dataset consists of 271K source lines of code (SLOC), 331K lines of JavaDoc, 1214K test suite SLOC and 223.5K test cases. Similar to the SF110 dataset, each input sample of this dataset contains the test prefix, MUT and docstring. We refer to this dataset as OracleEval25.
@@ II-B2 TOGLL
From the fine-tuning results, we identify CodeGen-350M and CodeParrot-110M as the top two models due to their high accuracy. We also select the three best-performing prompts: P4, P5, and P6. As different fine-tuned models may generalize differently to unseen data, and due to the differences between datasets, different prompts may work better. For this reason, we conduct a small-scale study, performing unseen inference with these two models and three prompts on projects from the OracleEval25 dataset. Our results suggest that even though CodeGen-350M is three times larger than CodeParrot-110M, CodeParrot-110M fine-tuned with P6 performs better on the unseen project than CodeGen on several metrics: total input processed, compilation error, and false positive count. Furthermore, the inference time with CodeGen-350M is longer than CodeParrot. Consequently, we define TOGLL as a fine-tuned CodeParrot-110M operating on P6. We perform rest of the study with TOGLL on the unseen OracleEval25 dataset II-B1.
II-B3 Baseline
@@
TOGA is a neural method for both assertion and exception oracle generation [17]. TOGA utilizes CodeBERT [37] model as backbone and fine-tuned for the oracle generation task. TOGA takes the test prefix generated by EvoSuite, its associated MUT and docstring (if available) and generates either an assertion or an exception oracle. An example of EvoSuite generated test cases with assertion and exception oracle is shown in Figure 2. The prefix part is marked with yellow color. The first test prefix involves pushing the number 10 onto the stack, followed by popping this value from the stack. For this scenario, an assertion oracle is anticipated to validate that the value retrieved is indeed 10. The second test attempts to pop from an empty stack, expecting the MUT to throw an exception. If no exception is thrown, the test should fail.
@@
We selected TOGA as our baseline method for several reasons: 1) TOGA represents the SOTA neural method, having surpassed specification, search, and neural-based techniques by detecting 57 bugs in Defects4J [36]; 2) TOGA utilizes the same set of information as P6, which is our most comprehensive prompt, thereby providing a fair basis for comparison; 3) TOGA has been previously evaluated on the OracleEval25 dataset, which is also employed in our study, ensuring consistency in our comparative analysis  [5].
II-B4 Metrics
With TOGLL, we generate test oracles for all projects from the OracleEval25 dataset. This dataset framework includes all necessary artifacts for executing the generated test oracles when integrated with the respective test prefixes. Consequently, we integrate the generated oracles into the test cases and execute the complete test suites. We calculate accuracy as the success rate, which denotes the percentage of non-empty test oracles that were successfully passed during test execution, indicating their alignment with expected program behavior. We compute the success rate using Equation 1
$\textrm{SuccessRate}(P)=\frac{T-(T_{ce}+T_{fp}+T_{em})}{T}$ (1)
where $T$ is the number of total test prefixes in project $P$, $T_{ce}$ is the number of compilation errors, $T_{fp}$ is the number false positive tests, and $T_{em}$ the number of empty oracles. We compute the SuccessRate metric for both assertion and exception oracles predicted by TOGLL and compare them to those of our baseline, TOGA.
@@ II-B4 Metrics
Figure 2: Test cases with assertion and exception oracles. The test prefix part is marked with yellow color.
@@
To assess the bug detection effectiveness of the TOGLL-generated oracles, we perform both mutation testing and real bug detection study on Defects4J. Mutation testing introduces small changes, or mutations, to the source code to create numerous slightly altered versions of the software, called mutants [38, 39]. Studies showed that there is a statistically significant correlation between mutant detection and real fault detection [40, 41, 42]. A test suite’s ability to detect these mutants can be a direct measure of its bug detection effectiveness [7, 6]. Therefore, we compute the mutant killing score as a metric to assess and compare the bug detection effectiveness of TOGLL-generated oracles with both EvoSuite and TOGA. Additionally, we also record unique mutants killing score which indicates the unique capability of each oracle generation method. Besides the mutation study, we also perform real bug detection study on the Defects4J dataset and compare the results with the TOGA baseline.
@@
TABLE I: Test oracle generation performance of LLMs on different prompts. Top two accuracies for each prompt are shown in bold. The last row shows prompt-wise average.
@@
III Experimental Study
To assess the efficacy of LLM-generated test oracles, this study investigates several key areas: effectiveness of different code LLMs in generating test oracles, the impact of different prompts on the oracle generation accuracy, the ability of fine-tuned models to generalize to unseen data, comparative accuracy among different oracle types, diversity of the generated oracles and the overall efficacy of these oracles in detecting bugs. To this end, we answer the following five research questions.
RQ4: How strong are TOGLL-generated assertions at identifying unique mutants? TOGLL generated oracles kill nearly 10 times more unique mutants than prior SOTA method and fall only 17% short of EvoSuite’s bug detection effectiveness.
RQ5: How strong are TOGLL-generated oracles at detecting real bugs from Defects4J? TOGLL-generated exception and explicit assertion oracles detected 106% more bugs compared to the prior SOTA method TOGA.